# Self-propulsion of floating ice blocks caused by melting in water

### v2

@@ V.1 Propulsion model for ice melting in saltwater
As discussed in the introduction and recent reviews, the melting of fresh water ice bodies including icebergs, ice shelves, sea ice floes in the ocean is a complex subject [11, 12, 13]. Several regimes have been identified to distinguish and describe ice melting in the ocean [11], depending on the ice surface orientation, the presence or absence of turbulence, the stratification of the salt concentration and the water temperature. The temperature of maximal density $T_{c}$ decreases with the salinity and is absent for a salinity above approximately $24$ g kg-1, suppressing reverse buoyancy at low temperatures [13]. Experimental studies have investigated the melting of ice block in a homogeneous saltwater bath in the absence of salt stratification at temperatures above $\,{}^{\circ}$[51, 25, 26, 52, 53, 17, 46]. According to Yamada, et al. [26], a double diffusive convection flow occurs near immersed ice blocks under those conditions, which is driven by the sinking cooled salt water below the block and rising fresh meltwater above. We note also a non-monotonic dependence of the melting rate with the salinity has been reported in experiments, with different behavior at the side and the base of the ice block [51].
@@
A schematic of the flows generated by melting and the cooling of the bath water based on observations is shown in Fig. 14. As in the freshwater case, we consider the inverted inclined ice surface with $x^{\prime}$ and $y^{\prime}$ as the coordinates along and perpendicular to the ice block with the origin located at the bottom of the ice block. A layer of pure meltwater rises up the block while ice melting cools the saltwater layer below. The cooling leads the saltwater in that layer to descent because of its increased density relative to the bathwater further below. Therefore, we assume that the melting velocity of the ice $v_{m}$ is determined by thermal convection. Because the ice is in contact with fresh meltwater, the ice melts at a the melting temperature for fresh water, i.e. $T_{i}=T_{m}=0\,^{\circ}$. As justified later in Section V.2, we assume that the meltwater layer thickness is small compared to the thermal boundary layer $\delta_{T}$, and its temperature remains close to $T_{m}=0\,^{\circ}$. As in Sec. IV.1, we calculate the thickness of the unstable boundary layer $\delta_{i}$, and derive,
@@
where $T_{max}(s_{A})$ is the temperature at which density of salt water at salinity $s_{A}$ is at its maximum over the range $[T_{i},T_{b}]$, and $\Delta\rho=\rho(T_{max}(s_{A}),s_{A})-\rho(T_{b},s_{A})$ is the density difference between salt water at $T_{max}$ and salt water at $T_{b}$. For $s_{A}\leq 18$ g kg-1, $T_{max}(s_{A})=T_{c}(s_{A})$ and $\delta_{T}$ is evaluated as in freshwater with Eq. (4). When $s_{A}>18$ g kg-1, $\rho(T,s_{A})$ decreases monotonically between $T_{i}$ and $T_{b}$. Then, we assume $T_{max}(s_{A})=T_{i}$, which gives $\delta_{T}=\delta_{i}$.
@@
A plot of $v_{m}$ versus $s_{A}$ for various $T_{b}$ between $1\,^{\circ}$ and $30\,^{\circ}$ is shown in Fig. 15(a). We find that $v_{m}$ increases with $s_{A}$, because the density contrast between warm and cold water is enhanced by the salinity.
Figure 15: (a) The melting velocity evaluated using Eq. (23) increases with bath salinity for $T_{b}$ between 0 and $30\,^{\circ}$. The color bar maps the line color to $T_{b}$. (b) The theoretical ice block velocity as a function of bath salinity at selected bath temperatures $T_{b}$ ($L=110$ mm and $\theta=23^{\circ}$).
We obtain $v_{p}$ by substituting in Eq. (23) in Eq. (14). As in freshwater, $U_{b}$ is then obtained from Eq. (17) by balancing the propulsion force with the inertial drag. Fig. 15(b) shows that $U_{b}$ increases with salinity and bath temperature. For example, going from $s_{A}=0$ to $s_{A}=40$ g kg-1 increases the predicted velocity by 25% at $20\,^{\circ}$. This is in clear contradiction with the experimental data, which suggest a significant decrease in $U_{b}$ with $s_{A}$. In fact one needs to take into account that as pure meltwater rises up along the block, it carries momentum in the opposite direction to the layer of cooled water. We calculate this contribution to the global momentum balance in the following.
@@ V.2 Melt layer dynamics
We can discuss the approximations made thus far using this estimate for the meltwater thickness and flow speeds before actually calculating the contribution of the meltwater layer to $U_{b}$. The velocity of the meltwater current averaged over the melt-layer thickness $\langle w\rangle_{y^{\prime}}=\frac{2}{3}\,K\,(h(x^{\prime}))^{2}$. We have $v_{m}\sim 40$ $\mu$m $s^{-1}$ at $20\,^{\circ}$ from Fig. 15(a), $\Delta\rho/\rho\approx 0.03$ between freshwater and saltwater with $s_{A}=35$ g kg-1 at $0\,^{\circ}$. Thus, for a block of length $L=20$ cm, $h(L)\approx 0.5$ mm, $w=\langle w\rangle_{y^{\prime}}(x^{\prime}=L)\approx 10$ mm $s^{-1}$ and a Reynolds number $Re=\dfrac{w\,h}{\nu}\approx 5$. The assumption that the flow is laminar therefore appears reasonable. However, the laminar meltwater flow hypothesis would not be valid for ice blocks that are a few meters in length, because $Re$ would be of order $100$. Further, the advection time scale $L/\langle w\rangle_{y^{\prime}}$ is of order $20$ s, while the salt diffusion time scale across the melt-layer can be estimated as $h^{2}/D\approx 250$ s, and the temperature diffusion time scale is $h^{2}/\kappa\approx 2.5$ s. These estimates are consistent with our assumption that salt does not have time to diffuse into the meltwater layer, which consequently remains as pure water along the length of the ice block. It can be noted that $h$ is smaller than $\delta_{T}$ by only a factor of 4 at $20^{\circ}$, which means that neglecting $h$ when calculating $\delta_{T}$ and assuming a uniform melt layer temperature at $T_{m}$, are rather crude approximations.
@@ V.3 Competition between the meltwater layer flow and the cooled layer flow
Fig. 16(a) shows plots of $F_{\mathrm{m}}$ and $F_{\mathrm{c}}$ for various bath temperatures. While $F_{\mathrm{c}}$ systematically increase with salinity over all $T_{b}$, $F_{\mathrm{m}}$ decreases and notably changes sign. It can be noted that $F_{\mathrm{c}}$ is greater than $F_{\mathrm{m}}$, leading to net propulsion consistent with our experimental observations. Only at very low $s_{A}$ over a range unexplored in our experiments does the temperature effect on density outweighs that of salinity.
@@
Figure 16: (a) The propulsion force in salt water for various bath temperature $T_{b}$. Continuous line, $F_{\mathrm{c}}$ contribution of the cooled layer in salt water according to Eq. (15). Dashed line, $F_{m}$ contribution of the melt layer, Eq. (29). (b) The ice propulsion velocity $U_{b}$ as a function of bath salinity for an ice block of length 21 cm and inclination $25^{\circ}$ for different bath temperatures $T_{b}$, taking into account the contribution of the melt layer. The theoretical estimate is obtained using Eq. (30).
Fig. 16(b) shows a plot of $U_{b}$ calculated with Eq. (30) as a function of salinity at various $T_{b}$, where it can be noted that $U_{b}$ is positive over a wide range of $T_{b}$ for sufficiently large $s_{A}$. While $U_{b}$ increases for sufficiently low temperatures, it decreases for $T_{b}$ above $20\,^{\circ}$. The decreasing trend is roughly consistent with our experiments performed in warm baths with $T_{b}\approx 23\,^{\circ}$. Although our model underestimates the observed decrease, this may be expected because of the number of simplifying assumptions made, including the rather strong assumption that the layer of meltwater neither mixes with the layer of cooled water below, nor exerts drag.
It is important to emphasize that our model is valid only when melting is driven by convective instability. It does not apply when the thermal boundary layer is stably stratified, which occurs when the bath temperature is below the critical temperature for a given salinity, i.e. $T_{b}<T_{c}(s_{A})$. Our model cannot thus predict the melting velocity and the block velocity for $T_{b}=1\,^{\circ}$ and $s_{A}\leq 13$ g kg-1 as shown in Fig. 16. We would expect that the layer of cooled water, flowing upward in the $x^{\prime}$-direction while remaining attached to the block, would generate a reactive force that propels the block in the opposite direction. Then, we would expect that the layer of cooled water, flowing upwards in the $x^{\prime}$ direction while remaining attached to the block, would propel the block in the other direction.
@@ V.4 First-order estimate of the terminal velocity of the ice block in salt water
To estimate the magnitude of propulsion velocity at ocean salinity, where the contribution from bath cooling dominates, we can apply the previous line of reasoning used in Sec. IV.3. Because the water density maxima is absent for salinity above $24$ g kg-1, we also do not consider the density contrast between $T_{c}$ and $T_{b}$, but rather between $T_{m}$ and $T_{b}$. Thus, Eq. (19) gives us,
@@
Because we assume that the ice remains in contact with the fresh meltwater layer, $T_{m}=0\,^{\circ}$, and not the lower value expected if the ice is in direct contact with saltwater. Then, for $T_{b}\approx 20\,^{\circ}$, we find $\Delta\rho/\rho\approx 3\times 10^{-3}$, $\delta_{T}\approx 1.3$ mm, and $U_{b,sc}\approx 8$ mm s-1, which overestimates $U_{b}$ by at least a factor 2 compared to experimental measurements. Nonetheless, Eq. (31) provides a correct order of magnitude for the block speed in saltwater. If we assume $T_{m}$ decreases with salinity, $U_{b,sc}\approx 8$ mm s-1 with $T_{m}=-1.90\,^{\circ}$ for a salinity of $35$ g kg-1, which is the same order of magnitude. Therefore, the decrease in the melting temperature with salinity has a little effect on the ice block velocity in warm water, if the ice melting is driven by thermal convection.
@@ VI Relevance of self-propulsion to natural icebergs
Icebergs found in the oceans are typically on the order of 100 meters in size, irregular shaped and composed of frozen freshwater, formed by the calving of ice sheets in Antarctica and Greenland [1]. The iceberg shapes below the waterline are not known well [55, 1], and aside from tabular icebergs, the emerged parts are often asymmetric due to processes involved their formation and deterioration [56]. Consequently, there is no strong reason to assume that the immersed shape of icebergs is symmetrical. Icebergs are predominantly reported in seawater with temperatures between 00 and $5\,^{\circ}$[57]. As such, our measurements are not conducted under typical conditions in which icebergs are found. However, icebergs have been reported drifting in warmer waters near Newfoundland ( around $45^{\circ}$ latitude) in spring, where the ocean temperature can reach $\,{}^{\circ}$, according to the Arctic Environmental Response Management Application (ERMA) of the National Oceanic and Atmospheric Administration (NOAA), (see https://erma.noaa.gov/arctic). Since icebergs can drift thousand of kilometers and reach these warm waters, the propulsion mechanism discussed in our study may be relevant to some iceberg scenarios. Using the scaling law for salt water given by Eq. (31), we estimate the propulsion speed for an iceberg of size about $200$ m melting in a seawater at $10\,^{\circ}$, and find $U_{b}\approx 0.05$ m s-1. While smaller, it is of the same order of magnitude as the typical drift velocities of icebergs, which is about $0.1$ m s-1 [58, 59]. This estimate suggests that while melting-induced propulsion alone cannot fully account for iceberg drift, it may represent a non-negligible contributing factor.
Moreover, even though the melting dynamics is not completely characterized for natural icebergs, the presence of currents generated by melting and asymmetric geometry is sufficient to generate a propulsion force, as was also noted in Ref. [9]. The concentrated rising currents of fresh meltwater near icebergs are well documented in field measurements [60, 61, 1, 47]. In a recent study [47], coupled water velocity and temperature measurements were performed near a vertical iceberg face approximately 10 m deep, melting in seawater at $\,{}^{\circ}$. At a depth of 6.5 m, a strong, rising meltwater plume was observed, with a velocity of approximately 5 cm s-1 extending 20 cm from the ice wall, in the absence of external currents. The velocity and width of a buoyant plume evolve in $d^{1/3}$ and $d$ respectively [62], where $d$ is the vertical distance to plume’s starting point, which is the bottom of the iceberg in this case. Using Eq. (15), we find therefore at first order $F_{c}/W\sim\rho_{b}\times\delta_{x}\times v_{p}^{2}\sim 1.9$ N m-1. For comparisons, a typical wind of velocity $u_{W}\approx 5$ m s-1 acting on a surface $W\times H_{e}$, with $H_{e}\approx 1$ m representing the emerged height, the force due to the wind is about $F_{w}/W\sim\rho_{air}\,\times H_{e}\times u_{W}^{2}\sim 25$ N m-1. In this estimate, the melting induced propulsion is about 7.5% that the one due to the wind. For a larger iceberg of draft and width $200$ m, a recent model [62] estimates the plume thickness and velocity to be about $10$ m and $0.06$ m s-1 for an ocean temperature of $1\,^{\circ}$. Similar orders of magnitudes are also obtained for plumes under ice shelves [50], which are the glaciers attached to Antarctica and Greenland and give birth to most of the icebergs. Then, using again Eq. (15), a first order estimation gives $F_{c}\sim 7\times 10^{3}$ N, whereas the contribution of the wind of $5$ m s-1 acting on a surface of order $10\times 200$ gives $F_{w}\sim 5\times 10^{4}$ N. In this case, the ratio of melting to the wind contribution to propulsion is about 14%.
These estimates suggest that while meltwater-driven propulsion is generally weaker than wind forcing, it may still contribute a small but non-negligible amount to iceberg motion. Importantly, because this effect is driven by solutal (compositional) convection, it remains effective even in cold water. The magnitude and direction of this force depend on the iceberg’s asymmetry and the inclination of its submerged surfaces. We also note that the convection flow generated by heat transfer into the surrounding salt water during melting has not been directly identified in field studies. It is difficult to separate this thermally driven flow from ambient ocean currents, as it lacks associated salinity gradients and may extend over length scales comparable to the iceberg itself. According to our estimates shown in Fig. 16, this thermal contribution could dominate at sea surface temperatures above approximately $5\,^{\circ}$.
@@ VII Discussion and Summary
In freshwater baths the propulsion is driven by the detached thermal convection flow, the mechanism is therefore valid for $T_{b}$ above the temperature $T_{c}=3.98\,^{\circ}$ where the water density maximum. As shown in Fig. 16, the model remains valid when ice melting occurs in saltwater and is driven by thermal convection, i.e. $T_{b}>T_{c}(s_{A})$. As $T_{c}(s_{A})$ decreases with $s_{A}$ the validity domain of the propulsion mechanism is extended to lower temperatures as long as $T_{b}>T_{m}(s_{A}=0)=0\,^{\circ}C$. However, further experimental tests are required of melting driven propulsion as bath temperatures approach the melting temperature.
@@ Appendix A Data sets
Data Set Image example $L_{h}\times W\times H$ (mm3) $L$ (mm) $\theta\,(^{\circ})$ Comments and Visualization $U_{b}$ mm s-1
G 27 runs [Uncaptioned image] Prescribed-size     $100\times 100\times 50$ 112 $24.3$ Custom made silicon molds. $10.4\leq T_{b}\leq 30\,^{\circ}$. Shadowgraph Avg. $2.3$ range $[0.4,4.0]$
R 12 runs [Uncaptioned image] Typical-size    $210\times 85\times 60$ Avg. $220$    range $[216,222]$ Avg. $12.5$    range $[8,15.8]$ Inclined silicon molds partially filled. Variable dimensions. $10.3\leq T_{b}\leq 29\,^{\circ}$. Shadowgraph. Avg. $3.1$ range $[0.9,4.3]$
D 4 runs [Uncaptioned image] Typical-size    $100\times 100\times 40$ Avg. $94$    range $[79,106]$ Avg. $15.2$    range $[12,17.2]$ Rectangular clear ice blocks cut in two along the diagonal using a hot wire. No trapped air. $T_{b}=21.4\,^{\circ}$ Shadowgraph Avg. $2.4$ range $[2.0,2.7]$
L 9 runs [Uncaptioned image] Typical-size    $165\times 125\times 65$ Avg. $179$    range $[176,180]$ Avg. $15.7$    range $[11.8,19.5]$ Rectangular ice blocks cut in two along the diagonal using a hot wire. Low content in air.$20.9\leq T_{b}\leq 22.1\,^{\circ}$. Shadowgraph Avg. $3.36$ range $[2.7,4.5]$
F 6 runs [Uncaptioned image] Prescribed-size     $100\times 100\times 40$ $100$ 00 Clear ice. Rectangular ice blocks. No trapped air. $20.4\leq T_{b}\leq 21.4\,^{\circ}$. Shadowgraph Avg. $0.3$ range $[0.1,0.43]$
T 47 runs [Uncaptioned image] Typical-size    $160\times 130\times 60$ Avg. $163$    range $[72,235]$ Avg. $19.5$    range $[6,36]$ Inclined silicone molds partially filled. Variable dimensions.$22\leq T_{b}\leq 24\,^{\circ}$. Top view Avg. $3.5$ range $[1.7,6.0]$
Table 1: List of various data sets for experiments performed in a fresh water bath.
Data Set Image example $L_{h}\times W\times H$ (mm3) $L$ (mm) $\theta\,(^{\circ})$ Comments and Visualization $U_{b}$ mm s-1
Gs 8 runs [Uncaptioned image] Prescribed-size     $100\times 100\times 50$ 112 $24.3$ Salinity $31$ g kg-1 Custom made silicon molds. $21.7\leq T_{b}\leq 24\,^{\circ}$. Shadowgraph Avg. $1.9$ range $[1.2,2.5]$
[…]
Table 2: List of various data sets for experiments performed in a salt water bath.

### v3

@@ V.1 Propulsion model for ice melting in saltwater
[…]
@@
A schematic of the flows generated by melting and the cooling of the bath water based on observations is shown in Fig. 14. As in the freshwater case, we consider the inverted inclined ice surface with $x^{\prime}$ and $y^{\prime}$ as the coordinates along and perpendicular to the ice block with the origin located at the bottom of the ice block. A layer of pure meltwater rises up the block while ice melting cools the saltwater layer below. The cooling leads the saltwater in that layer to descent because of its increased density relative to the bathwater further below. Therefore, we assume that the melting velocity of the ice $v_{m}$ is determined by thermal convection. Because the ice is in contact with fresh meltwater, the ice melts at a the melting temperature for fresh water, i.e. $T_{i}=T_{m}=0\,^{\circ}$C. As justified later in Section V.2, we assume that the meltwater layer thickness is small compared to the thermal boundary layer $\delta_{T}$, and its temperature remains close to $T_{m}=0\,^{\circ}$C. As in Sec. IV.1, we calculate the thickness of the unstable boundary layer $\delta_{i}$, and derive,
@@
where $T_{max}(s_{A})$ is the temperature at which density of salt water at salinity $s_{A}$ is at its maximum over the range $T_{i},T_{b}]$, and $\Delta\rho=\rho(T_{max}(s_{A}),s_{A})-\rho(T_{b},s_{A})$ is the density difference between salt water at $T_{max}$ and salt water at $T_{b}$. For $s_{A}\leq 18$ g kg-1, $T_{max}(s_{A})=T_{c}(s_{A})$ and $\delta_{T}$ is evaluated as in freshwater with Eq. (4). When $s_{A}>18$ g kg-1, $\rho(T,s_{A})$ decreases monotonically between $T_{i}$ and $T_{b}$. Then, we assume $T_{max}(s_{A})=T_{i}$, which gives $\delta_{T}=\delta_{i}$.
@@
A plot of $v_{m}$ versus $s_{A}$ for various $T_{b}$ between $1\,^{\circ}$C and $30\,^{\circ}$C is shown in Fig. 15(a). We find that $v_{m}$ increases with $s_{A}$, because the density contrast between warm and cold water is enhanced by the salinity.
Figure 15: Model predicting the melting velocity in salt water. (a) The melting velocity evaluated using Eq. (23) increases with bath salinity for $T_{b}$ between 0 and $30\,^{\circ}$C. The color bar maps the line color to $T_{b}$. (b) The theoretical ice block velocity as a function of bath salinity at selected bath temperatures $T_{b}$ ($L=110$ mm and $\theta=23^{\circ}$).
We obtain $v_{p}$ by substituting in Eq. (23) in Eq. (14). As in freshwater, $U_{b}$ is then obtained from Eq. (17) by balancing the propulsion force with the inertial drag. Fig. 15(b) shows that $U_{b}$ increases with salinity and bath temperature. For example, going from $s_{A}=0$ to $s_{A}=40$ g kg-1 increases the predicted velocity by 25% at $20\,^{\circ}$C. This is in clear contradiction with the experimental data, which suggest a significant decrease in $U_{b}$ with $s_{A}$. In fact one needs to take into account that as pure meltwater rises up along the block, it carries momentum in the opposite direction to the layer of cooled water. We calculate this contribution to the global momentum balance in the following.
@@ V.2 Melt layer dynamics
[…]
@@ V.3 Competition between the meltwater layer flow and the cooled layer flow
Fig. 16(a) shows plots of $F_{\mathrm{m}}$ and $F_{\mathrm{c}}$ for various bath temperatures. While $F_{\mathrm{c}}$ systematically increase with salinity over all $T_{b}$, $F_{\mathrm{m}}$ decreases and notably changes sign. It can be noted that $F_{\mathrm{c}}$ is greater than $F_{\mathrm{m}}$, leading to net propulsion consistent with our experimental observations. Only at very low $s_{A}$ over a range unexplored in our experiments does the Only at very low $s_{A}$ over a range unexplored in our experiments does the temperature effect on density outweighs that of salinity.
@@
Figure 16: Theoretical estimation of melting-driven propulsion forces and corresponding terminal velocities in salt water. (a) The propulsion force in salt water for various bath temperature $T_{b}$. Continuous line, $F_{\mathrm{c}}$ contribution of the cooled layer in salt water according to Eq. (15). Dashed line, $F_{m}$ contribution of the melt layer, Eq. (28). (b) The ice propulsion velocity $U_{b}$ as a function of bath salinity for an ice block of length 21 cm and inclination $25^{\circ}$ for different bath temperatures $T_{b}$, taking into account the contribution of the melt layer. The theoretical estimate is obtained using Eq. (29). The rescaled measured terminal velocity $\chi_{s}\,U_{b}$ for $T_{b}=23\,^{\circ}$ C from Fig. 12 are superimposed. These measurement points are approximately distributed around the theoretical curves for $15$, 2020, and $25\,^{\circ}$C. The experimental dispersion and uncertainty of measurements are too large for a quantitative comparison with the model.
Fig. 16(b) shows a plot of $U_{b}$ calculated with Eq. (29) as a function of salinity at various $T_{b}$, where it can be noted that $U_{b}$ is positive over a wide range of $T_{b}$ for sufficiently large $s_{A}$. While $U_{b}$ increases for sufficiently low temperatures, it decreases for $T_{b}$ above $20\,^{\circ}$C. Given the large experimental dispersion, our experiments performed in warm baths with $T_{b}\approx 23\,^{\circ}$C are consistent with the predicted trend. The experimental data points are spread around the theoretical curves for $T_{b}=15$, $T_{b}=20$ and $T_{b}=25\,^{\circ}$ C. We note also that our model relies on several assumptions made, including the rather strong assumption that the layer of meltwater neither mixes with the layer of cooled water below, nor exerts drag.
It is important to emphasize that our model is valid only when melting is driven by convective instability. It does not apply when the thermal boundary layer is stably stratified, which occurs when the bath temperature is below the critical temperature for a given salinity, i.e. $T_{b}<T_{c}(s_{A})$. Our model cannot thus predict the melting velocity and the block velocity for $T_{b}=1\,^{\circ}$C and $s_{A}\leq 13$ g kg-1 as shown in Fig. 16. We would expect that the layer of cooled water, flowing upward in the $x^{\prime}$-direction while remaining attached to the block, would generate a reactive force that propels the block in the opposite direction. Then, we would anticipate that the layer of cooled water, flowing upwards in the $x^{\prime}$ direction while remaining attached to the block, would propel the block in the other direction.
@@ V.4 First-order estimate of the terminal velocity of the ice block in salt water
To estimate the magnitude of propulsion velocity at ocean salinity, where the contribution from bath cooling dominates, we can apply the previous line of reasoning used in Sec. IV.3. Because the water density maxima is absent for salinity above $24$ g kg-1, we also do not consider the density contrast between the temperature of maximum density $T_{c}$ and bath temperature $T_{b}$, but rather between melting temperature $T_{m}$ and bath temperature $T_{b}$. Thus, Eq. (19) gives us,
@@
Because we assume that the ice remains in contact with the fresh meltwater layer, $T_{m}=0\,^{\circ}$C, and not the lower value expected if the ice is in direct contact with saltwater. Then, for $T_{b}\approx 20\,^{\circ}$C, we find $\Delta\rho/\rho\approx 3\times 10^{-3}$, $\delta_{T}\approx 1.3$ mm, and $U_{b,sc}\approx 8$ mm s-1, which overestimates $U_{b}$ by at least a factor 2 compared to experimental measurements. Nonetheless, Eq. (30) provides a correct order of magnitude for the block speed in saltwater. If we assume $T_{m}$ decreases with salinity, $U_{b,sc}\approx 8$ mm s-1 with $T_{m}=-1.90\,^{\circ}$C for a salinity of $35$ g kg-1, which is the same order of magnitude. Therefore, the decrease in the melting temperature with salinity has a little effect on the ice block velocity in warm water, if the ice melting is driven by thermal convection.
@@ VI Relevance of self-propulsion to natural icebergs
[…]
These estimates suggest that while meltwater-driven propulsion is generally weaker than wind forcing, it may still contribute a small but non-negligible amount to iceberg motion. Importantly, because this effect is driven by solutal (compositional) convection, it remains effective even in cold water. The magnitude and direction of this force depend on the iceberg’s asymmetry and the inclination of its submerged surfaces. We also note that the convection flow generated by heat transfer into the surrounding salt water during melting has not been directly identified in field studies. It is difficult to separate this thermally driven flow from ambient ocean currents, as it lacks associated salinity gradients and may extend over length scales comparable to the iceberg itself. According to our estimates shown in Fig. 16, this thermal contribution could dominate at sea surface temperatures above approximately $5\,^{\circ}$C.
@@ VII Discussion and Summary
In freshwater baths the propulsion is driven by the detached thermal convection flow, the mechanism is therefore valid for $T_{b}$ above the temperature $T_{c}=3.98\,^{\circ}$C where the water density maximum. As shown in Fig. 16, the model remains valid when ice melting occurs in saltwater and is driven by thermal convection, i.e. $T_{b}>T_{c}(s_{A})$. As $T_{c}(s_{A})$ decreases with $s_{A}$ the validity domain of the propulsion mechanism is extended to lower temperatures as long as $T_{b}>T_{m}(s_{A}=0)=0\,^{\circ}C$. However, further experimental tests are required of melting driven propulsion as bath temperatures approach the melting temperature.
@@ Appendix A Data sets
See pages - of TableI.pdf
See pages - of TableII.pdf
See pages - of TableIII.pdf